How much heat is required to raise an 820g copper pan from 12?C to 88?C if (a) the pan is empty; (b) the pan contains 1.1kg of water; and (c) the pan contains 4.5kg of mercury?

?Qa =? kcal

?Qb =? kcal

?Qc =? kcal
Answer #1

specific heat of copper = 0.385 J/g C

So a) Qa=820*0.385*76=23993.2 Joules

b) Heat reguired to raise both water and copper temperature

so Qb=23993.2+4.18*1100*76=373441.2 Joules

c) heat required to raise mercury Temperature along with copper so

Qc=23993.2+(0.14*4500*76)=71873.2 Joules
Answer #2

1) Heat required = 1 cal/gC x 0.385 Joules x 820g x (88-12) C= 23.9 KJ

2) Heat required = 1 cal/gC x (0.385+4.18) Joules x 820g x (88-12) C= 284.5 KJ

3) Heat required = 1 cal/gC x (0.385+4.18+0.14) Joules x 820g x (88-12) C= 293.2 KJ